# Model-Based Transfer Learning for Contextual Reinforcement Learning

### v1

@@ Abstract
Deep reinforcement learning is a powerful approach to complex decision making. However, one issue that limits its practical application is its brittleness, sometimes failing to train in the presence of small changes in the environment. This work is motivated by the empirical observation that directly applying an already trained model to a related task often works remarkably well, also called zero-shot transfer. We take this practical trick one step further to consider how to systematically select good tasks to train, maximizing overall performance across a range of tasks. Given the high cost of training, it is critical to choose a small set of training tasks. The key idea behind our approach is to explicitly model the performance loss (generalization gap) incurred by transferring a trained model. We hence introduce Model-Based Transfer Learning (MBTL) for solving contextual RL problems. In this work, we model the performance loss as a simple linear function of task context similarity. Furthermore, we leverage Bayesian optimization techniques to efficiently model and estimate the unknown training performance of the task space. We theoretically show that the method exhibits regret that is sublinear in the number of training tasks and discuss conditions to further tighten regret bounds. We experimentally validate our methods using urban traffic and standard control benchmarks. Despite the conceptual simplicity, the experimental results suggest that MBTL can achieve greater performance than strong baselines, including exhaustive training on all tasks, multi-task training, and random selection of training tasks. This work lays the foundations for investigating explicit modeling of generalization, thereby enabling principled yet effective methods for contextual RL.
@@ 1 Introduction
Deep reinforcement learning (DRL) has made remarkable strides in addressing complex problems across various domains [24, 31, 1, 4, 10, 12, 23]. Despite these successes, DRL models often exhibit brittleness when exposed to small variations in task settings, significantly limiting their scalability and generalizability [28, 40]. Such variations in Markov Decision Processes (MDPs) can be conceptualized as contextual Markov Decision Processes (CMDP), where slight differences in MDP definitions lead to multiple, closely related but different environments [14, 27]. Traditional DRL approaches typically require independent training for each new task variant, a process that is both resource-intensive and impractical for real-time applications. Multi-task training is a relatively inexpensive way to train but still has limited model capacity issues. In light of this, there is a pressing need for more robust approaches that enhance generalization across different tasks. It is crucial to differentiate between training robustness, which seeks resilience of training across a variety of tasks, and model robustness, which focuses on the resistance of a model to environmental perturbations. Contextual reinforcement learning (CRL) offers a framework for modeling changes in tasks [6]. In our exploration of applying DRL to solve multiple related tasks, we identify computational challenges, particularly the high costs and complexities associated with training separate models for each task. While techniques like CARL [6] or multitask RL [36, 2] address these challenges, they still remain computationally expensive and inefficient when applied across a broad context range.
@@
Figure 1: Generalization gap in Cartpole CMDPs. The solid lines illustrate the actual generalized performance after zero-shot transfer across contexts, with the source context indicated by a dotted line.
Figure 2: Model-Based Transfer Learning (MBTL) for solving contextual MDPs. MBTL framework strategically selects optimal tasks to train by modeling the generalization gap that arises when transferring a model to different tasks. For each target task, we deploy the most effective trained model, indicated by solid arrows. We evaluate our framework on a suite of continuous and discrete tasks, including standard control benchmarks [6] and urban traffic benchmarks [34, 19].
@@
We build upon zero-shot transfer, which has shown potential in directly applying a policy trained in one MDP (source task) to another (target task) without adaptation. This often leads to a performance degradation known as the ‘generalization gap’ [16, 20]. For instance, Figure 1 exemplifies the concept of the generalization gap within a Cartpole CMDP, a range of scenarios in which the agent must balance a pole on a moving cart with a different mass of pole. The performance degrades as the target task diverges from the source task, illustrating an increasing generalization gap. Given that training is expensive yet zero-shot transfer is cheap, we are interested in optimally selecting a set of source (training) tasks, such that performance on the target range of tasks is maximized. In this article, we devise strategies to incorporate the structure of the generalization gap to better estimate the value of training. Specifically, we investigate how explicit modeling of the generalization gap structure lends itself to principled and effective algorithms for the selection of training tasks within CMDP.
Hence, this article introduces Model-Based Transfer Learning (MBTL), a novel framework for studying generalization in reinforcement learning (Figure 2). We model the explicit form of the generalization gap in zero-shot transfer and use it in estimating how valuable selecting the source task is. Given a small number of samples in earlier steps, we leverage Bayesian optimization methods to estimate training performance with limited information, thus significantly reducing the necessity for exhaustive training [35, 37].
This paper contributes to the field of reinforcement learning by introducing MBTL framework for efficient source task selection, aiming to solve multiple related tasks with enhanced transferability and fewer required models. Theoretical analyses detail the sublinear cumulative regret of these methods and provide conditions for achieving tighter regret bounds. We validate our methods in simulated traffic management scenarios and standard control benchmarks. Our findings provide foundations for investigating explicit modeling of generalization, enabling efficient and practical model-based solutions for solving multiple related tasks for robust training in DRL.
@@ 2 Preliminaries and notation
We leverage the framework of CMDPs [14, 27, 6] to formulate the problem of solving multiple related tasks using reinforcement learning.
Contextual MDP (CMDP). A standard MDP is defined by the tuple $M=(S,A,P,R,\rho)$ where $S$ represents the state space, $A$ is the action space, $P$ denotes the transition dynamics, $R$ is the reward function, and $\rho$ is the distribution over initial states. In a CMDP, denoted by $c\mathcal{M}(x)=(S,A,P_{x},R_{x},\rho_{x})$, these components are parameterized by a context variable $x\in X$, which can influence both dynamics and rewards [14, 27]. Consider zero-shot transfer from a source task (MDP) to solve another target task (MDP) in the CMDP. The notation $x$ is used to represent a context for the source task, with each task $x$ belonging to the finite set of CMDPs denoted as $X$. The context of the target task is similarly denoted by $x^{\prime}$ and is also an element of $X$. In our framework, the expected return of a MDP $c\mathcal{M}(x)$ with context $x$ is denoted by $J(x)$. We differentiate between estimated values $\hat{J}(x)$ and observed outcomes $J(x)$, with the latter measured after training and evaluation.
Generalization gap via zero-shot transfer. Zero-shot transfer involves applying a model trained on a source task $c\mathcal{M}(x)$ to a different target task $c\mathcal{M}(x^{\prime})$, with $x^{\prime}$ also belonging to the set $X$. We observe the generalized performance, denoted by $U(x^{\prime};x)$, by evaluating the target task $x^{\prime}$ based on the model trained using source task $x$ via zero-shot generalization. As the concept of “transferability” has been discussed in literature [32, 3], we define the generalization gap as the absolute performance difference in average reward when transferring from source task $x$ to target task $x^{\prime}$:
$\underbrace{\Delta J(x,x^{\prime})}_{\text{Generalization gap}}=\underbrace{U(x^{\prime};x)}_{\text{Generalized performance}}-\underbrace{J(x)}_{\text{Source task performance}}.$ (1)
@@ 3 Problem formulation
Sequential source task selection problem. The selection of source tasks in CMDPs is key to solving the overall CMDP [3]. Specifically, we consider the sequential source task selection (SSTS) problem, which seeks to maximize the expected performance across a dynamically selected set of tasks. This problem is framed as a sequential decision problem, in which the selection of tasks is informed through feedback from the observed task performance of the tasks selected and trained thus far. The notation $x_{k}$ indicates the selected source task at the $k$-th transfer step, where $k$ ranges from 1 to $K$.
@@
The sequential source task selection (SSTS) problem seeks to optimize the overall expected performance across a CMDP $c\mathcal{M}(\cdot)$ by selecting one task $x\in X$ at each stage. Formally, at each selection step $k$, we aim to choose a task $x_{k}$ such that the cumulative estimated performance is maximized:
$\max_{x_{k}}\ \ V(x_{k};x_{1},...,x_{k-1})\quad\text{s.t. }x_{k}\in X\setminus\{x_{1},...,x_{k-1}\}.$ (2)
where $V(\cdot)$ is defined in the following. Intuitively, we can keep track of the highest performance achievable by any model trained thus far. Formally, after we train the $k^{\text{th}}$ model at source task $x_{k}$, we update the generalized performance, denoted by $U(x^{\prime};x_{(1:k)})$, by comparing the generalized performance from the model trained using task $x_{k}$. We recursively define $U$ based on the previous observations, as follows:
@@
$U(x^{\prime};x_{(1:k)})=\max\left(U(x^{\prime};x_{k}),U(x^{\prime};x_{(1:k-1)})\right)\quad\forall x^{\prime}\in X\quad\text{if }k>1.$ (3)
For simplicity, we denote the sequence $x_{1},x_{2},...,x_{k}$ by $x_{(1:k)}$. We can calculate the observed expected generalized performance, denoted as $V(x_{(1:k)})=\mathbb{E}_{x^{\prime}\in X}\left[U(x^{\prime};x_{(1:k)})\right]$.
@@
The state at each step $k$ is defined by the best performance for each task, achieved by models trained in earlier stages, represented as $U(x^{\prime};x_{(1:k-1)})$ for each target task $x^{\prime}$. The action at each step involves choosing a new task $x_{k}$, aimed at optimizing generalized performance based on current knowledge. The cumulative reward is quantified by the observed performance of the combined task sequence, $V(x_{(1:k)})$, providing a direct measure of the strategic value of each decision. In general, SSTS exhibits stochastic transitions, for example due to randomness in RL training. For simplicity, in this work, we assume deterministic transitions; that is, training MDP $x$ will always yield the same performance $J(x)$ and generalization gap $\Delta J(x,x^{\prime}),\forall x^{\prime}\in X$. The problem’s horizon is defined by $|X|$, indicating that all models will be trained unless a specific termination condition is met. This termination condition is determined by reaching a predefined level of suboptimality.
Model-based transfer learning (MBTL). All methods that solve SSTS are transfer learning methods, by definition. We additionally designate a method that solves SSTS to be model-based if an explicit structure is imposed upon $\Delta J(\cdot,\cdot)$. To the best of the authors’ knowledge, this is the first work to study such methods in the context of reinforcement learning.
4 Methods for Model-Based Transfer Learning
Algorithm 1 Model-based Transfer Learning (MBTL)
0:CMDPs $c\mathcal{M}(x)$, Task $x\in X$, Transfer budget $K$
0:$\boldsymbol{\pi}$ and $V$ Initialize : $J,V=0\ \forall x\in X$, $\boldsymbol{\pi}=\{\}$, $k=1$
1:while $k\leq K$ do
2:    $x_{k}\leftarrow\textbf{NextTask}(J_{k-1},V_{k-1})$
3:    $\pi_{k}\leftarrow\text{Train}(\mathcal{M}(x_{k}))$
4:    $\boldsymbol{\pi}\leftarrow\boldsymbol{\pi}\cup\{\pi_{k}\}$
5:    Update $J(x),V(x)$
6:    $k\leftarrow k+1$
7:end while
8:return  $\boldsymbol{\pi}$ and $V$
(a) Pseudo-greedy strategy
(b) Equidistant strategy
Figure 3: Naive methods for solving SSTS. Note that the normalized performance of the trained tasks is assumed to always be 1 (known and constant).

### v2

@@ Abstract
Deep reinforcement learning (RL) is a powerful approach to complex decision making. However, one issue that limits its practical application is its brittleness, sometimes failing to train in the presence of small changes in the environment. Motivated by the success of zero-shot transfer—where pre-trained models perform well on related tasks—we consider the problem of selecting a good set of training tasks to maximize generalization performance across a range of tasks. Given the high cost of training, it is critical to select training tasks strategically, but not well understood how to do so. We hence introduce Model-Based Transfer Learning (MBTL), which layers on top of existing RL methods to effectively solve contextual RL problems. MBTL models the generalization performance in two parts: 1) the performance set point, modeled using Gaussian processes, and 2) performance loss (generalization gap), modeled as a linear function of contextual similarity. MBTL combines these two pieces of information within a Bayesian optimization (BO) framework to strategically select training tasks. We show theoretically that the method exhibits sublinear regret in the number of training tasks and discuss conditions to further tighten regret bounds. We experimentally validate our methods using urban traffic and standard continuous control benchmarks. The experimental results suggest that MBTL can achieve up to 50x improved sample efficiency compared with canonical independent training and multi-task training. Further experiments demonstrate the efficacy of BO and the insensitivity to the underlying RL algorithm and hyperparameters. This work lays the foundations for investigating explicit modeling of generalization, thereby enabling principled yet effective methods for contextual RL. Code is available at https://github.com/jhoon-cho/MBTL/.
@@ 1 Introduction
Deep reinforcement learning (DRL) has made remarkable strides in addressing complex problems across various domains [29, 36, 1, 4, 11, 13, 27]. Despite these successes, DRL algorithms often exhibit brittleness when exposed to small variations like different number of lanes, weather conditions, or flow density in traffic benchmarks [20], significantly limiting their scalability and generalizability [19]. Such variations can be modeled using the framework of contextual Markov Decision Processes (CMDP), where task variations can be parameterized within a context space [15, 32, 6].
There are two predominant solution modalities for CMDPs [22]: independent training and multi-task training. Independent training constructs a separate model for each task variant (say, $N\gg 1$), which is compute-intensive. At the other extreme, multi-task training constructs a single “universal” policy, and thus can be more compute-efficient, but suffers from model capacity and negative transfer issues [21, 44, 2, 41]. There is thus a need for more reliable training methodologies for generalization across tasks variants. In this work, we consider training an intermediate set of $K$ models, where $N>K>1$, in an effort to balance performance and efficiency; we refer to this strategy as multi-policy training.
A note on terminology. For brevity, we refer to task variants as tasks in the remainder of this article. We also use the language of task and context interchangeably. We emphasize that this work focuses on within-domain generalization (e.g., traffic signal control for intersection scenario variants) rather than across-domain generalization (e.g., distinct traffic control tasks). Additionally, it is crucial to differentiate between training reliability, which concerns the ability to reliably train models across tasks, and model reliability (or robustness), which concerns the resistance of a trained model to differences in tasks. This article is concerned with training reliability.
@@
Figure 1: Example generalization gap depicted for Cartpole CMDP. The solid lines show the true zero-shot transfer generalization performance across contexts. Source tasks are indicated by dotted lines.
@@
We build upon zero-shot transfer, a widely-used practical technique which directly applies a policy trained in one context (source task) to another (target task) without adaptation. This experiences performance degradation, also called generalization gap [17, 22]. For instance, Figure 1 depicts that the performance degrades as the target task diverges from the source task, corresponding to an increasing generalization gap. Nonetheless, leveraging the notion that training is expensive but zero-shot transfer is cheap, we are interested in optimally selecting a set of source (training) tasks, such that the generalization performance on the target range of tasks is maximized. In this article, we strategically select source tasks by explicitly modeling the generalization performance to estimate the value of training a new source task.
The main contributions of this work are:
  • We introduce Model-Based Transfer Learning (MBTL), a novel framework for studying generalization in reinforcement learning (Figure 2). To the best of our knowledge, this is the first work to explicitly model generalization performance for contextual RL. As such, our work opens the door for further investigation into reliable model-based methodologies for contextual RL.
  • We provide theoretical analysis for the sublinear regret of MBTL and give conditions for achieving tighter regret bounds.
  • We empirically validate our methods in urban traffic and standard continuous control benchmarks for contextual RL, observing up to 50x improvements in sample efficiency. We further include ablations on the components of the algorithm.
Figure 2: Overview illustration for Model-based Transfer Learning. (a) Gaussian process regression is used to estimate the training performance across tasks using existing policies; (b) marginal generalization performance (red area) is calculated using upper confidence bound of estimated training performance, generalization gap, and generalization performance; (c) selects the next training task that maximizes the acquisition function (marginal generalization performance); (d) once the selected task is trained, calculate generalization performance using zero-shot transfer.
The remainder of the paper is organized as follows. After introducing notation in Section 2, we formally define the problem in Section 3. A key contribution of our work is the introduction of a Gaussian process model acquisition function specifically tailored to the source task selection problem, which is detailed in Section 4. In Section 4.3, we provide a theoretical analysis of the regret bounds of our method, followed by an empirical evaluation across diverse applications in Section 5.
@@ 2 Preliminaries and notation
Contextual MDP. A standard MDP is defined by the tuple ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}M}=(S,A,P,R,\rho)$ where $S$ represents the state space, $A$ is the action space, $P$ denotes the transition dynamics, $R$ is the reward function, and $\rho$ is the distribution over initial states [43]. A contextual MDP (CMDP), denoted by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}=}(S,A,P_{x},R_{x},\rho_{x})_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x\in X}}$, is a collection of context-MDPs ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}_{x}}$ parameterized by a context variable $x$ within a context space $X$ (assumed bounded). The context variable $x$ can influence dynamics, rewards, and initial state distributions [15, 32, 6]. We define source task performance $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0};\text{Alg}})$ as follows: we train a policy $\pi_{x}$ on a task with the context $x\in X$ using RL algorithm Alg (e.g., PPO, SAC) and evaluate the policy by the expected return in the same MDP $\mathcal{M}_{x}$ with context $x$. For brevity, we will write it as $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x)$. We distinguish between estimated values $\hat{J}(x)$ and observed outcomes $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x)$, with the latter measured after training and evaluation.
Generalization gap via zero-shot transfer. Consider zero-shot transfer from the trained policy $\pi_{x}$ from a source task (context-MDP) to solve another target task (context-MDP) with the context $x^{\prime}\in X$ in the CMDP. Zero-shot transfer involves applying a policy trained on a source task ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}_{x}}$ to a different target task ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{M}_{x^{\prime}}}$, with $x,x^{\prime}\in X$. We observe the generalization performance, denoted by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{x},x^{\prime}})$, by evaluating the target task $x^{\prime}$ based on the policy trained using source task $x$ via zero-shot generalization. We define the generalization gap as the absolute performance difference in average reward when transferring from source task $x$ to target task $x^{\prime}$:
$\underbrace{\Delta J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})}_{\text{Generalization gap}}=\underbrace{J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x)}_{\text{Source task performance}}-\underbrace{J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}})}_{\text{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}Generalization} performance}}.$ (1)
@@ 3 Problem formulation
Sequential source task selection problem. The selection of source MDPs from the CMDPs is key to solving the overall CMDP [3]. We therefore introduce the sequential source task selection (SSTS) problem, which seeks to maximize the expected performance across a dynamically selected set of tasks. This problem is cast as a sequential decision problem, in which the selection of tasks is informed through feedback from the observed task performance of the tasks selected and trained thus far. The notation $x_{k}$ indicates the selected source task at the $k$-th transfer step, where $k$ ranges from 1 to $K$. For brevity, we will denote $\pi_{x_{k}}$ as $\pi_{k}$. We denote the sequence $x_{1},x_{2},...,x_{k}$ by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x_{1:k}}$ and $\pi_{1},\pi_{2},...,\pi_{k}$ by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k}}$.
@@
[…]
@@
${J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x_{k}},x^{\prime};\pi_{1:k-1}})$ $=\max\left({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}J(\pi_{k},x^{\prime})},{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1},x^{\prime}})\right)\quad\forall x^{\prime}\in X\quad\text{if }k>1.$ (2)
And ${J}(\pi_{1:1},x)\equiv{J}(\pi_{1},x)$. Then, the overall sequential decision problem can be written as:
$\max_{x_{k}}\ \ {V}(x_{k};{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1}})=\max_{x_{k}}\mathbb{E}_{x^{\prime}\sim\mathcal{U}(X)}\left[{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x_{k}},x^{\prime};\pi_{1:k-1}})\right]\quad\text{s.t. }x_{k}\in X\setminus x_{1:k-1}.$ (3)
@@
The state at each step $k$ is defined by the best generalization performance for each task, achieved by policies trained in earlier stages, represented as $J({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{1:k-1},x^{\prime}})$ for each target task $x^{\prime}$. The action at each step is choosing a new task $x_{k}$. In general, SSTS exhibits stochastic transitions, for example due to randomness in RL training. For simplicity, in this work, we assume deterministic transitions; that is, training context-MDP $x$ will always yield the same performance ${J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x}},x)$ and generalization gap $\Delta{J}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\pi_{x},x^{\prime}}),\forall x^{\prime}\in X$. The problem’s maximum horizon is defined by $|X|$, but can be terminated early if conditions are met (e.g., performance level, training budget).
4 Model-Based Transfer Learning (MBTL)
In this section, we introduce an algorithm called Model-based Transfer Learning to solve the SSTS problem. MBTL models the generalization performance in two parts: 1) the performance set point, modeled using Gaussian processes, and 2) performance loss (generalization gap), modeled as a linear function of contextual similarity. MBTL combines these two pieces of information within a Bayesian optimization (BO) framework to sequentially select training tasks that maximize generalization performance. The high-level pseudo-code of the MBTL algorithm is given below.
MBTL Pseudo-code 1:Input: Task set, Gaussian Process (GP) model, and training budget 2:while training budget remains do 3:    Select next source task $\leftarrow$ Task with the highest acquisition value based on estimated training performance, generalization gap, and generalization performance, determined by the current GP model and tailored acquisition function 4:    Train a policy on the selected source task using reinforcement learning, then evaluate its training performance 5:    Update the GP model with the selected source task training performance to estimate the training performance of tasks in the task set 6:end while 7:Output: Set of policies that maximizes generalized performance across task set